OpenStudy (anonymous):

Use F=ma to explain why more massive objects accelerate at the same rate as less massive objects? More massive objects fall faster because gravity pulls on them harder. In F=ma, the force of gravity is the same for all objects. More massive objects require more force to accelerate at the same rate as less massive objects. In F=ma, if the acceleration is the same, but the mass goes up, then the force has to go up too Less massive objects must fall slower because in F=ma, air resistance is a smaller force that makes things fall together.

OpenStudy (anonymous):

the last option is: Gravity is the same for all objects so F=ma does not matter. i believe it is the last option. is that correct?

OpenStudy (anonymous):

oh no, i take it back. the second option maybe?

OpenStudy (emmatassone):

The correct answer is the second one, because all the objects have the same acceleration when they fall in the earth aproximately

OpenStudy (anonymous):

thank you :) could i ask another question? this is the last one.

OpenStudy (emmatassone):

sure

OpenStudy (anonymous):

The formula for KE = ½mv2. The formula for PE = mgh. If the mass of an object doubles, then the KE of that object should ________. double triple quadruple not change

OpenStudy (anonymous):

would it double ?

OpenStudy (emmatassone):

\[K_i= \frac{ 1 }{ 2 }m_Iv ^{2}\]now if \[m_f=2m_i \rightarrow K_f=\frac{ 1 }{2 }m_fv ^{2}= \frac{ 1 }{ 2 }(2m_i)v ^{2}=2(\frac{ 1 }{ 2 }m_{i}v ^{2})\rightarrow 2K_i\]

OpenStudy (emmatassone):

so it would be the double

OpenStudy (anonymous):

ahhh thank you!!

OpenStudy (anonymous):

i get it now!